### v2

@@ Abstract
An important challenge in machine learning is to predict the initial conditions under which a given neural network will be trainable. We present a method for predicting the trainable regime in parameter space for deep feedforward neural networks, based on reconstructing the input from subsequent activation layers via a cascade of single-layer auxiliary networks. For both the MNIST and CIFAR10 datasets, we show that a single epoch of training of the shallow cascade networks is sufficient to predict the trainability of the deep feedforward network, thereby providing a significant reduction in overall training time. We achieve this by computing the relative entropy between reconstructed images and the original inputs, and show that this probe of information loss is sensitive to the phase behaviour of the network. Moreover, our approach illustrates the network’s decision making process by displaying the changes performed on the input data at each layer. Our results provide a concrete link between the flow of information and the trainability of deep neural networks, further explaining the role of criticality in these systems.
@@ 1 Introduction
Despite the impressive empirical success of deep neural networks (DNNs) the theoretical principles underlying deep learning remain relatively poorly understood. However, concepts and ideas from physics have proven increasingly useful for improving both our theoretical understanding and the practical use of DNNs. The relationship between neural networks and statistical mechanics in particular has a rich history (see Bahri-review for a review, as well as references therein), and connections between deep learning and the renormalization group (RG) have been explored in, among others, mehta2014exact ; Koch-Janus_2017 ; LenggenhagerOT ; Erdmenger:2021sot ; Kline_2022 ; Berman:2022uov ; JRAZ . More recently, various correspondences between DNNs and quantum field theory have been developed (see Halverson:2020trp ; Grosvenor:2021eol and subsequent works) and techniques from QFT have been applied more generally at both a structural Roberts:2021fes ; Berman:2024pax ; Berman:2022mak and dynamical Cohen level.
An important issue for machine learning is to determine for which initial configurations neural networks are optimally trainable. Current state-of-the-art approaches aim at optimizing these configurations by grid searches over parameter space that involve training multiple network configurations (see e.g. optuna_2019 ). These grid searches may however represent an obstacle to the cost effectiveness of neural networks. It is therefore highly desirable to find the optimal initial network configuration prior to performing any training. Here we achieve this by presenting an efficient framework, based on a physics-inspired concept that we name reconstruction entropy, that predicts the trainable network configurations. In our experiments for the MNIST and CIFAR10 datasets, we find a speed-up of up to two orders of magnitude. Moreover, our approach allows to identify which features of the data stored in the hidden layers are particularly important for the network’s decision-making process.
@@
The starting point for the present work is the observation that many deep neural networks exhibit a phase transition between an ordered and chaotic phase poole2016exponential ; schoenholz2017deep ; xiao2018dynamical ; chen2018dynamical . These are characterized by vanishing and exploding gradients, respectively, and hence networks initialized in either phase are difficult or impossible to train. Intuitively, this is because information about the input data cannot propagate through the network to affect the cost function, since correlations either die-out in the ordered (low-temperature) phase, or are washed-out by noise in the chaotic (high-temperature) phase. However, the boundary between these phases is described by a continuous phase transition at which the correlation length diverges, allowing theoretically infinite signal propagation, and hence trainability of arbitrarily deep networks. In other words, a precondition for trainability is that the depth of the network must not exceed the scale set by the correlation length. Thus, increasingly deep networks must be initialized closer to this so-called “edge of chaos” in order to be trainable. Determining the location of the edge of chaos in phase space is thus of significant practical benefit, since it avoids the need for computationally costly grid searches over the space of parameters. [^1] Here we mean the weights and biases that parametrize the phase space. The selection of appropriate hyperparameters such as batch size, learning rate, etc., are of secondary importance in the sense that no arrangement of hyperparameters can redeem a network in the untrainable regime. Conversely, if the network does not appear to train at some arbitrary point in phase space, it is a priori unclear whether the selection of parameters or hyperparameters is the problem. The present work aims to resolve this. For related work on trainability vs. generality, see e.g. Xiao2020 and references therein.
@@
Initial works poole2016exponential ; schoenholz2017deep on deep feedforward networks determined the location of the critical line in the idealized regime in which the number neurons per layer $N$ is taken to infinity. [^2] This framework also holds if one allows different numbers of neurons $N_{\ell}$ in different layers $\ell$, provided all $N_{\ell}\to\infty$ at the same rate. Informally, all one requires is that the distribution of each layer is approximately Gaussian. Due to the central limit theorem, the distribution of neurons in each layer becomes normal as $N\to\infty$, and hence the network becomes a Gaussian process (see rasmussen2005gaussian for a comprehensive treatment). However, as real-world networks are necessarily of finite width, the theoretical prediction for the critical point as well as the correlation length itself differs from empirical results; this can be seen already in the initial works above, and was shown even more starkly in Erdmenger:2021sot . While significant mathematical machinery has been brought to bear on the problem of predicting finite-width corrections to this so-called mean field [^3] Most works in the machine learning literature refer to the infinite width limit as mean field theory. As discussed in Grosvenor:2021eol however, this is not quite right for two reasons: conceptually, there is no continuum limit and hence no bona fide fields in these previous works; and technically, the mean field theory is also not necessarily the same as the result from the central limit theorem. Since the latter is ultimately all that is required, we refer to this line of work exclusively under this banner. result Grosvenor:2021eol ; Naveh_2021 ; bordelon2023dynamics ; PhysRevLett.120.248301 ; Li_2020 , accurately determining the trainable regime for a given network architecture remains an open challenge.
@@
In light of the above, it is pertinent to ask whether other tools from physics may be applied to determine the location of the edge of chaos prior to training (that is, sparing the need for computationally costly grid searches over the entire phase space). Ultimately, the problem is to determine the conditions under which information, suitably quantified, can propagate through the entire network. In Erdmenger:2021sot , parallels between subsequent layers in a DNN and successive steps of real-space RG were found by analyzing the information flow through the network via the Kullback-Leibler (KL) divergence or relative entropy,
@@
The activations of each layer $\ell$ are then fed as inputs to a shallow auxiliary network, which is trained to reconstruct the original preactivations $\mathbf{z}^{\ell-1}$. Note that the training of the auxiliary networks is done sequentially: that is, at $\ell\!=\!1$, a shallow cascade network $\mathcal{C}^{1}$ is trained to reconstruct the original input $\mathbf{z}^{0}=\mathbf{x}$ from the activations $\mathbf{z}^{1}$. At $\ell\!=\!2$, a second shallow network is trained to reconstruct $\mathbf{z}^{1}$ from $\mathbf{z}^{2}$. Once trained, this layer is combined with the previous auxiliary layer to form cascade $\mathcal{C}^{2}$ (see fig. 1), so that a reconstructed image can then be obtained from $\mathbf{z}^{2}$, and so on. Thus, the auxiliary cascade network $\mathcal{C}^{\ell}$ consists of $\ell$ individually trained layers, each of which acts as a shallow decoder for the corresponding layer in the feedforward network (henceforth referred to as “the DNN”). In this sense, the relationship between the feedforward network and the cascades resembles that of an autoencoder, though in this case the activations of each individual layer in the DNN are taken as a sequence of latent spaces. Importantly, since the cascades are trained as a collection of shallow networks, training is extremely quick, and can be completed with high accuracy within a single epoch. Since the components (layers) of the cascades are trained individually, they are always in the trainable regime for any reasonable choice of parameter values. [^6] It is possible for networks to be untrainable even at criticality if the activation function saturates Bukva:2023ksv . Furthermore, the relative entropy computed between the reconstructed image from cascade $\mathcal{C}^{\ell}$ and the original input provides a more indicative measure of meaningful information loss, as this minimizes artifacts associated with qualitatively identical content being encoded in different ways. [^7] We note that computing the entropy without using the cascades, i.e., with the layers of the feedforward network itself, merely results in noise. This is illustrated in fig. 1 by the images reconstructed from deeper and deeper layers of the DNN becoming increasingly blurred as information is lost that the cascades are unable to reconstruct. To again employ an analogy with autoencoders, we do not want to compute the relative entropy between the data and the latent representation, but between the data and its reconstruction. We note that this method also avoids the constraint, previously imposed in Erdmenger:2021sot , that the widths of all layers of the feedforward network be identical to avoid issues with normalizing the distributions, since here we are always computing the relative entropy between vectors in the same space. In the subsequent, when using the term ‘relative entropy’, we always refer to the relative entropy between the reconstruction and the input.
@@
Note that we have used $\rho(x)$ to denote the (continuous) probability distribution function appearing in this expression, to avoid confusion with the (discrete) probability mass function $p(x)$ appearing in (1). We construct $\rho(x)$ as follows (see fig. 2). For fixed $\ell$, we consider a set of reconstructed images obtained for a random selection of inputs. For all images in this set, choose the same pixel, denoted by its location $(i,j)$ (with $i,j\in\{1,\ldots,\sqrt{N}\}$ where, e.g., for MNIST $N\!=\!784$), and normalize so that $\rho_{ij}(x)$ is the distribution of values taken by pixel $(i,j)$ over the set of reconstructed images. If this set is sufficiently large, then $\rho$ becomes approximately Gaussian, and therefore the entropy (3) is simply
@@
where $\sigma_{ij}$ is the standard deviation of $\rho_{ij}(x)$. We define the reconstruction differential entropy as the sum of the differential entropies for all pixels in the set of reconstructed images, $S\coloneqq\frac{1}{N}\sum_{ij}S_{ij}$. For a random collection of inputs, differences between pixel values between images will result in a relatively high variance. As shown in fig. 2 however, as the information is lost in successive layers, the reconstructed images will tend to converge, resulting in a small ($\sigma_{ij}^{2}\ll 1$) variance and hence an increasingly negative $S$. The use of a large sample set in this method reduces the sensitivity to fluctuations, whereas the method described above using (1), while providing a more interpretable representation of information flow, can by construction only compare a single reconstruction image with the original input. In the results section below, we use both the relative entropy (1) and the differential entropy (3) within the reconstruction approach. While the differential entropy shows greater numerical stability, we emphasize that the results of both methods for computing the entropy associated with reconstructed images yields qualitatively and quantitatively similar results for the cutoff $\ell^{*}$.
@@
A further advantage of the reconstruction method is that the encoding of the information in the hidden layers, after reconstruction, is similar to the encoding of the information in the input layer. This renders the information stored in the hidden layers accessible to the observer and provides an approach to understanding the decision-making process of neural networks. An example for this is given in fig. 1, where the digit seven can be identified in all reconstructions. It is possible to identify certain features of the reconstructed images as particularly relevant for the network’s decision-making. This may be of particular use when using machine learning in daily-life situations, where it is crucial to be able to understand the decisions taken by the network and to verify their suitability for the given task.
@@
Figure 2: Illustration of the two methods for computing the reconstruction entropy. Panel (a) shows the reconstructed image from layers $\ell=0,3,5$ (left to right) for a trainable but untrained DNN. Panels (b) and (c) show the same for a network in the untrainable regime with two different inputs. One sees in the third column ($\ell\!=\!5$) that for untrainable networks, different inputs converge to a common reconstruction as information is lost. The red and blue boxes around images in panels (b) and (c) correspond to the computation of the differential and relative entropy, respectively, which are illustrated in more detail in panels (d) and (e). For the differential entropy, we construct a probability distribution function for each individual pixel, drawn from the set of reconstruction images. (Panel (d) illustrates this for a sample size of 3; in practice, we use a sample size of 100.) For the relative entropy, we compare sequential reconstruction images with the corresponding input; note that the variation between individual pixels is also important in this approach, but as only two images are compared at a time, the probability mass function is discrete (represented by histograms in panel (e)). In the bottom panel (f), we show the relative entropy for an untrained network in the stable, critical, and chaotic regimes. The stable and chaotic regimes are characterized by saturation of the relative entropy after a certain point we call the cutoff. For networks at criticality, there is no such cutoff (by default, it is equal to the total depth $L$). This provides the basis for our prediction of trainability without needing to train the feedforward network; we only need to train the series of shallow networks (i.e., the individual layers) that comprise the cascades, cf. fig. 1.

### v3

@@ Abstract
An important challenge in machine learning is to predict the initial conditions under which a given neural network will be trainable. We present a method for predicting the trainable regime in parameter space for deep feedforward neural networks (DNNs) based on reconstructing the input from subsequent activation layers via a cascade of single-layer auxiliary networks. We show that a single epoch of training of the shallow cascade networks is sufficient to predict the trainability of the deep feedforward network on a range of datasets (MNIST, CIFAR10, FashionMNIST, and white noise), thereby providing a significant reduction in overall training time. We achieve this by computing the relative entropy between reconstructed images and the original inputs, and show that this probe of information loss is sensitive to the phase behaviour of the network. We further demonstrate that this method generalizes to residual neural networks (ResNets) and convolutional neural networks (CNNs). Moreover, our method illustrates the network’s decision making process by displaying the changes performed on the input data at each layer, which we demonstrate for both a DNN trained on MNIST and the vgg16 CNN trained on the ImageNet dataset. Our results provide a technique for significantly accelerating the training of large neural networks.
@@ 1 Introduction
Despite the impressive empirical success of deep neural networks (DNNs), the theoretical principles underlying deep learning remain relatively poorly understood. A particularly important challenge for machine learning is to determine for which initial configurations neural networks are optimally trainable. Current state-of-the-art approaches aim at optimizing these configurations by grid searches over parameter space that involve training multiple network configurations (see e.g. optuna_2019 ). These grid searches may however represent an obstacle to the cost effectiveness of neural networks, and are not even feasible for larger networks used in industry. To reduce this computational burden, significant theoretical effort has been devoted to determine the initial conditions under which a given network will be trainable. Here, we present an alternative approach that significantly accelerates the process of selecting suitable parameters by predicting trainable regions in phase space without any training of the DNN.
Related work. The vast majority of existing theoretical and experimental approaches take place in the infinite-width approximation, in which the number of neurons $N$ per layer is taken to infinity. [^1] This framework also holds if one allows different numbers of neurons $N_{\ell}$ in different layers $\ell$, provided all $N_{\ell}\to\infty$ at the same rate. Informally, all one requires is that the distribution of each layer is approximately Gaussian. Due to the central limit theorem, the distribution of neurons in each layer becomes normal as $N\to\infty$, and hence the network becomes a Gaussian process (see rasmussen2005gaussian for a comprehensive treatment). This mathematical simplification allows many results to be derived exactly.
One important line of work in this direction centers on the neural tangent kernel (NTK) Jacot2018 . In the infinite-width limit, the NTK becomes frozen, obtaining a duality between training the network under gradient descent and kernel methods, allowing one to study the training dynamics analytically in this regime, see e.g. Nakazato2024AST ; Zhao2021ZerOII . However, results based on the constancy of the NTK become less accurate if the ratio of depth to width is finite, as is the case in real-world networks. For example, of particular relevance to the present work is Seleznova2022 , which examines how the behaviour of the NTK changes in different regions of phase space (we shall elaborate on this momentarily) when both width and depth are taken to infinity. More generally however, networks in the infinite-width NTK parametrization do not actually learn features Yang2020FeatureLI , which further underscores the need for alternative approaches.
Another important approach to the problem of fine-tuning large (industry-scale) networks is hyperparameter transfer: the basic idea is to find the optimum initialization for a small neural network, and then deduce how that parametrization will scale to a large one. This approach has been spearheaded by Yang and collaborators Yang2022TensorPV (see alsoMei2018AMF ; Chizat2018OnTG for previous work), who demonstrated that so-called maximal-update parametrization of shallow networks can be successfully transfered to larger models. This has since been generalized to deeper networks in Yang2023TensorPV ; Bordelon2023DepthwiseHT . Notably however, as commented in sec. 1.3 of Yang2022TensorPV , this approach breaks down when the block depth of the ResNet is greater than 1. Additionally, as will be seen below, this approach is still much more computationally intensive than the one we present here, since one still has to find the optimal conditions for, and successfully train, the smaller network; for the case studied in Yang2022TensorPV for example, this required 5000 training steps, whereas our approach allows one to determine this in a single epoch.
A particularly relevant application of trainability prediction is to signal propagation in ResNets. For example, Yang2017MeanFR showed that the optimal intialization variances for ResNets may depend on the depth, which poses a challenge for existing theoretical approaches (cf. Yang2022TensorPV above). Related work Li2021TheFI ; Li2022TheNC seeks to make progress by again taking both the width $N$ and depth $L$ to infinity, so that the parameter $L/N$ remains finite, as mentioned above. Interestingly, these limits appear to commute in the special case of ResNets Hayou2023WidthAD . In general, this is not the case, since the non-linearities accumulate with depth, causing a loss of theoretical control. At a technical level, this is because the finite-width corrections to the Gaussian Process approximation can be written as a parametric expansion in $L/N$ Roberts:2021fes ; Grosvenor:2021eol . Nonetheless, while idealized, this limit remains an important tool for studying these networks due to the afforded level of theoretical control.
To summarize these developments, it is highly desirable – and even necessary, in the case of larger networks used in industry – to find the optimal initial network configuration prior to performing training. Here, we present a highly efficient alternative to the approaches discussed above. In our experiments for the MNIST and CIFAR10 datasets for example, we find a speed-up of up to two orders of magnitude. Moreover, our approach allows to identify which features of the data stored in the hidden layers are particularly important for the network’s decision-making process. While this work is primarily a proof of concept, we have also generalized our method to CNNs and ResNets, and tested it on the additional datasets FashionMNIST and ImageNet.
@@
Our approach is based on a physics-inspired concept that we name reconstruction entropy, and builds on the rapidly growing intersection of physics and machine learning in the past few years, wherein concepts and ideas from physics have proven increasingly useful for improving both our theoretical understanding and the practical use of DNNs. The relationship between neural networks and statistical mechanics in particular has a rich history (see Bahri-review for a review, as well as references therein), and connections between deep learning and the renormalization group (RG) have been explored in, among others, mehta2014exact ; Koch-Janus_2017 ; LenggenhagerOT ; Erdmenger:2021sot ; Kline_2022 ; Berman:2022uov ; Howard:2024bbi . More recently, various correspondences between DNNs and quantum field theory have been developed (see Halverson:2020trp ; Grosvenor:2021eol and subsequent works), providing a rigorous explanatory framework for the finite-width effects discussed above, and techniques from QFT have been applied more generally at both a structural Roberts:2021fes ; Berman:2024pax ; Berman:2022mak and dynamical Cohen ; Howard:2024bbi level.
@@
More specifically, the starting point for the present work is the observation that, as alluded above, many deep neural networks exhibit a phase transition between an ordered and chaotic phase poole2016exponential ; schoenholz2017deep ; xiao2018dynamical ; chen2018dynamical . These are characterized by vanishing and exploding gradients, respectively, and hence networks initialized in either phase are difficult or impossible to train. Intuitively, this is because information about the input data cannot propagate through the network to affect the cost function, since correlations either die-out in the ordered (low-temperature) phase, or are washed-out by noise in the chaotic (high-temperature) phase. However, the boundary between these phases is described by a continuous phase transition at which the correlation length diverges, allowing theoretically infinite signal propagation, and hence trainability of arbitrarily deep networks. In other words, a precondition for trainability is that the depth of the network must not exceed the scale set by the correlation length. Thus, increasingly deep networks must be initialized closer to this so-called “edge of chaos” in order to be trainable. Determining the location of the edge of chaos in phase space is thus of significant practical benefit, since it avoids the need for computationally costly grid searches over the space of parameters. [^2] Here we mean the weights and biases that parametrize the phase space. The selection of appropriate hyperparameters such as batch size, learning rate, etc., are of secondary importance in the sense that no arrangement of hyperparameters can redeem a network in the untrainable regime. Conversely, if the network does not appear to train at some arbitrary point in phase space, it is a priori unclear whether the selection of parameters or hyperparameters is the problem. The present work aims to resolve this. For related work on trainability vs. generality, see e.g. Xiao2020 and references therein.
@@
Initial works poole2016exponential ; schoenholz2017deep on deep feedforward networks determined the location of the critical line in above-mentioned idealized regime in which the width of the network is taken to infinity. However, as real-world networks are necessarily of finite width, the theoretical prediction for the critical point as well as the correlation length itself differs from empirical results; this can be seen already in the initial works above, and was shown even more starkly in Erdmenger:2021sot . While significant mathematical machinery has been brought to bear on the problem of predicting finite-width corrections to this central limit theorem [^3] Most works in the machine learning literature refer to the infinite width limit as mean field theory. As discussed in Grosvenor:2021eol however, this is not quite right for two reasons: conceptually, there is no continuum limit and hence no bona fide fields in these previous works; and technically, the mean field theory is also not necessarily the same as the result from the central limit theorem. Since the latter is ultimately all that is required, we refer to this line of work exclusively under this banner. result Grosvenor:2021eol ; Naveh_2021 ; bordelon2023dynamics ; PhysRevLett.120.248301 ; Li_2020 , accurately determining the trainable regime for a given network architecture remains an open challenge.
@@
In light of the above difficulties, it is pertinent to ask whether other tools from physics may be applied to determine the location of the edge of chaos prior to training (that is, sparing the need for computationally costly grid searches over the entire phase space). Ultimately, the problem is to determine the conditions under which information, suitably quantified, can propagate through the entire network. In Erdmenger:2021sot , parallels between subsequent layers in a DNN and successive steps of real-space RG were found by analyzing the information flow through the network via the Kullback-Leibler (KL) divergence or relative entropy,
@@
The activations of each layer $\ell$ are then fed as inputs to a shallow auxiliary network, which is trained to reconstruct the original preactivations $\mathbf{z}^{\ell-1}$. Note that the training of the auxiliary networks is done sequentially: that is, at $\ell\!=\!1$, a shallow cascade network $\mathcal{C}^{1}$ is trained to reconstruct the original input $\mathbf{z}^{0}=\mathbf{x}$ from the activations $\mathbf{z}^{1}$. At $\ell\!=\!2$, a second shallow network is trained to reconstruct $\mathbf{z}^{1}$ from $\mathbf{z}^{2}$. Once trained, this layer is combined with the previous auxiliary layer to form cascade $\mathcal{C}^{2}$ (see fig. 1), so that a reconstructed image can then be obtained from $\mathbf{z}^{2}$, and so on. Thus, the auxiliary cascade network $\mathcal{C}^{\ell}$ consists of $\ell$ individually trained layers, each of which acts as a shallow decoder for the corresponding layer in the feedforward network (henceforth referred to as “the DNN”). In this sense, the relationship between the feedforward network and the cascades resembles that of an autoencoder, though in this case the activations of each individual layer in the DNN are taken as a sequence of latent spaces. Importantly, since the cascades are trained as a collection of shallow networks, training is extremely quick, and can be completed with high accuracy within a single epoch. Since the components (layers) of the cascades are trained individually, they are always in the trainable regime for any reasonable choice of parameter values. [^6] It is possible for networks to be untrainable even at criticality if the activation function saturates Bukva:2023ksv . Furthermore, the relative entropy computed between the reconstructed image from cascade $\mathcal{C}^{\ell}$ and the original input provides a more indicative measure of meaningful information loss, as this minimizes artifacts associated with qualitatively identical content being encoded in different ways. [^7] We note that computing the entropy without using the cascades, i.e., with the layers of the feedforward network itself, merely results in noise. This is illustrated in fig. 1 by the images reconstructed from deeper and deeper layers of the DNN becoming increasingly blurred as information is lost that the cascades are unable to reconstruct. To again employ an analogy with autoencoders, we do not want to compute the relative entropy between the data and the latent representation, but between the data and its reconstruction. We note that this method also avoids the constraint, previously imposed in Erdmenger:2021sot , that the widths of all layers of the feedforward network be identical to avoid issues with normalizing the distributions, since here we are always computing the relative entropy between vectors in the same space. In what follows, when using the term ‘relative entropy’, we always refer to the relative entropy between the reconstruction and the input. [^8] As an aside, we note that the Kullback-Leibler divergence has also been employed to improve trainability in Peer , though this approach contrasts strongly with ours. There, a new, entropy-inspired term is added to the loss function, which contains another hyperparameter that must be obtained via an additional grid search. In contrast, we do not alter the loss function; furthermore, as emphasized below, the novel auxiliary cascades introduced in our approach significantly reduce the need for such hyperparameter searches.
@@
Note that we have used $\rho(x)$ to denote the (continuous) probability distribution function appearing in this expression, to avoid confusion with the (discrete) probability mass function $p(x)$ appearing in (1). We construct $\rho(x)$ as follows (see fig. 2). For fixed $\ell$, we consider a set of reconstructed images obtained for a random selection of inputs. For all images in this set, choose the same pixel, denoted by its location $(i,j)$ (with $i,j\in\{1,\ldots,\sqrt{N}\}$ where, e.g., for MNIST $N\!=\!784$), and normalize so that $\rho_{ij}(x)$ is the distribution of values taken by pixel $(i,j)$ over the set of reconstructed images. If this set is sufficiently large, then $\rho$ becomes approximately Gaussian [^11] We emphasize that we are not referring to the distribution of pixels within a given image such as discussed in, e.g., Hyvrinen2009NaturalIS ., and therefore the entropy (3) is simply
@@
where $\sigma_{ij}$ is the standard deviation of $\rho_{ij}(x)$. Empirical support for this statement is given in appendix A.5. We define the reconstruction differential entropy as the sum of the differential entropies for all pixels in the set of reconstructed images, $S\coloneqq\frac{1}{N}\sum_{ij}S_{ij}$. For a random collection of inputs, differences between pixel values between images will result in a relatively high variance. As shown in fig. 2 however, as the information is lost in successive layers, the reconstructed images will tend to converge, resulting in a small ($\sigma_{ij}^{2}\ll 1$) variance and hence an increasingly negative $S$. The use of a large sample set in this method reduces the sensitivity to fluctuations, whereas the method described above using (1), while providing a more interpretable representation of information flow, can by construction only compare a single reconstruction image with the original input. In the results section below, we use both the relative entropy (1) and the differential entropy (3) within the reconstruction approach. While the differential entropy shows greater numerical stability, we emphasize that the results of both methods for computing the entropy associated with reconstructed images yields qualitatively and quantitatively similar results for the cutoff $\ell^{*}$.
@@
A further advantage of the reconstruction method is that the encoding of the information in the hidden layers, after reconstruction, is similar to the encoding of the information in the input layer. This renders the information stored in the hidden layers accessible to the observer and provides an approach to understanding the decision-making process of neural networks. An example for this is given in fig. 1, where the digit seven can be identified in all reconstructions. A more involved example is given in fig. 7 using a colored figure of two animals and the vgg16 convolutional neural network simonyan2015deepconvolutionalnetworkslargescale . It is possible to identify certain features of the reconstructed images as particularly relevant for the network’s decision-making. This may be of particular use when using machine learning in daily-life situations, where it is crucial to be able to understand the decisions taken by the network and to verify their suitability for the given task.
@@
Figure 2: Illustration of the two methods for computing the reconstruction entropy. Panel (a) shows the reconstructed image from layers $\ell=0,3,5$ (left to right) for a trainable but untrained DNN. Panels (b) and (c) show the same for a network in the untrainable regime with two different inputs. One sees in the third column ($\ell\!=\!5$) that for untrainable networks, different inputs converge to a common reconstruction as information is lost. The red and blue boxes around images in panels (b) and (c) correspond to the computation of the differential and relative entropy, respectively, which are illustrated in more detail in panels (d) and (e). For the differential entropy, we construct a probability distribution function for each individual pixel, drawn from the set of reconstruction images. (Panel (d) illustrates this for a sample size of 3; in practice, we use a sample size of 100.) For the relative entropy, we compare sequential reconstruction images with the corresponding input; note that the variation between individual pixels is also important in this approach, but as only two images are compared at a time, the probability mass function is discrete. The normalized activation for each pixel is represented by histograms in panel (e)). In the bottom panel (f), we show the relative entropy for an untrained network in the stable, critical, and chaotic regimes. The stable and chaotic regimes are characterized by saturation of the relative entropy after a certain point we call the cutoff. For networks at criticality, there is no such cutoff (by default, it is equal to the total depth $L$). This provides the basis for our prediction of trainability without needing to train the feedforward network; we only need to train the series of shallow networks (i.e., the individual layers) that comprise the cascades, cf. fig. 1.